OpenStudy (anonymous):

An airplane us asked to stay within a holding pattern near ab airport. The function d(x)=90sin(0.87x)+160 represents the distance d, in miles, that the airplane is from the airport at time "x", in minutes. During the first 17 minutes after the plane enters the holding pattern at what time "x" is the plane exactly 120 miles from the airport?

OpenStudy (anonymous):

can i help plz

OpenStudy (anonymous):

solve \[70\sin(0.65t) + 150 =100\]

OpenStudy (anonymous):

subtract 150 and get \[70\sin(0.65t) =-50 \] divide by 70 to get \[\sin(0.65t)=-\frac{5}{7}\]

OpenStudy (anonymous):

then take the arcsine of both sides, get \[0.65t=\sin^{-1}(\frac{5}{7})\] you will need a calculator

OpenStudy (anonymous):

then divide both sides by \(.65\)